he is 21 years old

Found 2 solutions by scott8148, magnifik0:
Answer by scott8148(5890)

X-10= age of eric 10 years ago
Eric is now 3 times as old as he was 10 years ago: 3(X-10)
X is his age now

3(X-10)=X
3X-30=X
2X=30
X=15

Answer by magnifik0(1)

he is 21 years old